What are the steps to solve this question:
Find the sum of the first n terms of each arithmetic sequence.
-5, -1, 3....; n=17

Answer by jim_thompson5910(35256)

In this case, the first term is a1 = -5 and the common difference is d = 4 (eg: -1 minus -5 = 4)
-------------------------------------------------------
Sn = n*(a1 + an)/2

Sn = n*(a1 + a1 + d*(n-1))/2

Sn = n*(2a1 + d*(n-1))/2

Sn = n*(2*(-5) + 4*(n-1))/2

S17 = 17*(2*(-5) + 4*(17-1))/2

S17 = 459
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------